An astronaut on the moon throws a baseball upward. The astronaut is 6​ ft, 6 in.​ tall, and the initial velocity of the ball is 30 ft per sec. The height s of the ball in feet is given by the equation s equals -2.7t^2 + 30t + 6.5​, where t is the number of seconds after the ball was thrown. Complete parts a and b.
a. After how many seconds is the ball 12 ft above the​ moon's surface?

Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52778)

.

Solve the quadratic equation


    -2.7t^2 + 30t + 6.5 = 12


using the quadratic formula.


The two solutions will give you two time moments.


Answer by Alan3354(69443)

s equals -2.7t^2 + 30t + 6.5​, where t is the number of seconds after the ball was thrown. Complete parts a and b.
-------------------
a. After how many seconds is the ball 12 ft above the​ moon's surface?
s(t) = -2.7t^2 + 30t + 6.5 = 12
2.7t^2 - 30t + 5.5 = 0
Solve for t
---
The smaller value is ascending.
The larger is descending.